SHOW ALL WORK. Preferred if you showed how to do it on both financial calculator and by hand.

1. You plan to retire in 35 years. At the end of each year, you plan on saving $15,000, and your bank pays you 2% annual interest. How much will you have saved by the time you retire?

A. $749,917

B. $374,979

C. $525,000

D. $262, 514

2.  A bond with a face value of $1,000 pays a 10% (APR) semiannual coupon, and matures in 10 years. Similar bonds trade at a YTM of 8% (APR). What is the price of the bond?

A. $1,000

B. $679.52

C. $1,134.20

D. $1,135. 91

PROBLEMS 3 and 4 ARE CONNECTED

3. A major car manufacturing firm issues a 20 year $1,000,000 bond at par. The bond pays a 6% (APR) semiannual coupon. 5 years later, the Federal Reserve Board cuts the fed funds rate, causing interest rates for similar firms to fall to 4% (APR). What is the new bond price?

A. $1,222,368

B. $1,223,965

C. $1,000,000

D. $1,273,555

4.  . If you bought the bond at issue and held it to maturity, what is the effective annual rate (EAR) that you earned?

A. 6%

B. 6.09%

C. 4%

D. 4.04%

5.  The nominal rate quoted in the market is 8% per annum, and interest is paid every month. What is the effective annual rate (EAR)?

A. 8%

B. 8.3%

C. 0.67%

D. 96%
Answer #1

(1) Investment Tenure = 35 years, Annual Savings = $ 15000, Annual Interest = 2%

Final Value of Savings at retirement = 15000 x (1.02)^(34) + 15000 x (1.02)^(33) + ............+ 15000 x (1.02) + 15000 = [(1.02)^(35) - 1] / [1.02 - 1] x 15000 = $ 749917.2 ~ $ 749917

Hence, the correct option is (A)

The problem can be solved using the BA II Plus calculator as described below:

- Input PV = $ 0, N = 35, I/Y = 2, PMT = $ 15000, CMPT - > FV = $ 749917

NOTE: PLease raise separate queries for solutions to the remaining unrelated questions as one query is restricted to the solution of only one complete question with a maximum of four sub-parts.